Solve the system of equations by graphing
x+y=8
x-y=-6
What is the solution of the system of equations?

Answer by vleith(2983)

See this video --> http://collabvsl.wetpaint.com/page/Solving+Systems+of+Equations
The solution in this case is the point that is the intersection of the two given lines.
There are several methods to solve this.
We will use addition
x + y = 8
x - y = -6 add these two
-------------
2x = 2
x = 1
If x =1, then y is
x + y = 8
1 + y = 8
y = 7
So the solution is the point (1,7)
Check your answer using the second eqaution
x - y = -6
does 1 - 7 = -6 ??? yes it does !